OpenStudy (anonymous):

Solve the following using the quadratic formula. Round to the nearest tenth, if necessary. 2x^2 + 9x + 4 = 0

OpenStudy (anonymous):

D= b ^2 - 4ac D= 9*9 -4 *2*4=81-64=17 x = (-9(+or-)root17)/4